Showing posts with label betting theories. Show all posts
Showing posts with label betting theories. Show all posts
Thursday, October 9, 2008
1-3-2-6 BETTING SYSTEM
The name of this system says it all. It is similar to the Paroli system. It is based on the premise that you can win four times in a row. Your initial bet is 1 unit, the second 3 units, the third 2 units and the fourth 6 units. Let’s assume that each unit is $10 and the odds are 1:1 - even money. The first bet is $10. When winning, $10 is added to the $20 on the table making the second bet $30. When winning again on the second bet, there would be $60 on the table. Of this you take down $40 and the third bet is now $20. If the third bet wins, you will have $40 on the table to which you add $20 making a total of $60 for the fourth bet. If the fourth bet wins, there would be a total of $120 on the table, all of which is net profit. Now all the bet is taken down and you start the system all over again at $10. Also, each time you lose, at any level, you start all over again at $10. If you lose the first bet, your loss is $10. The second level loss is $20 (because you added another $10). At the third level, a loss will give you a net profit of $20 (because you have taken down $40). At the fourth level, a loss leaves you breaking even (because you put back $20 out of the $40 taken down). The attraction of this system is that you risk $20 at a chance of making $120 net profit. This means you can lose six times at the worst level (second bet), and with one win (a set of four wins in a row) get your money back
Етикети:
betting theories
FLAT BETTING SYSTEM
Betting the same fixed amount on each play is known as Flat Betting. With this simple strategy you establish a bet size, say $100, and you bet this same amount each time whether you win or lose. Of course this flat bet size could be too large or too small in proportion to your bankroll if your stake grows from $1000 to $5000, or shrinks back down to $500! On the upside, the Flat Bet is very conservative, although its fixed amount bet constrains bankroll growth and makes it difficult to have a really outstanding year. On the downside, a severe string of losses can completely decimate a Flat Betting bankroll as the bet size becomes increasingly over-sized relative to a shrinking bankroll! It is widely believed that professional gamblers bet flat. The reasons for their choice might be the following:
Stakes are adjusted, if adjusted at all, according to odds (probabilities), not the recent history.
Stakes are usually small comparing to the gambling fund, thus the mental pressure is not so tough, and the prospect of bankruptcy is not so preeminent.
The profit shifts nice and easy, roller coaster ride excluded. Many take this to be a very healthy feature.
Ideal for flat beting are spreads and unders/overs in US Sports.
Етикети:
betting theories
Martingdale strategy
The Martingdale” is (in)famous betting strategy known as well as red&black roulette system. This system helps you rather control your stakes than choosing games.Theoretically it will gain you profit in sports betting independent of you ability to choose right game, but the risks to go broke are pretty high. Whe you loosefirst stake you wil adjust amount of next stake with a help of the table I will publish later. You will not just simply double amount but the rule is lower next game odds higher you stake must be. For example you lost 10 uit and next game/ticket odds are 2.00, your stake will be again 10,if you win you gained your lossess back,if you loose and next odds are again 2.00 you should bet 400 and not 800 as by simply doubling stake. It will help you got further with your money.
Етикети:
betting theories
Fibonacci sequence
Fibonacci sequence - betting system.
1-2-3-5-8-13-21-34. Go up one step with each loss, down 2 steps with each win. Note that every win pays for the two losses before it.
Example:
1. Bet $1 until you lose. Then bet $2.
2. If you win at $2, then return to step 1. If you lose, then bet $3.
3. If you win at $3, then return to step 1. If you lose, then bet $5.
4. If you win at $5, then return to step 2. If you lose, then bet $8.
5. If you win at $8, then return to step 3. If you lose, then bet $13.
6. If you win at $13, then return to step 4. If you lose, then bet $21.
7. If you win at $21, then return to step 5. If you lose, then bet $34.
8. If you win at $34, then return to step 6. If you lose, return to step 1.
Recommended variant for the high roller: instead of going down 2 steps with each win, go down 1 step with each win, until you win 2 in a row-then go down 2 steps. This variant may be as likely as any system to help you in the long run, and meanwhile you get some tremendous wins and excitement. There are numerous other variants. However the usual Fibonacci is extremely unsatisfactory for the high roller-you lose often and never win big. Whereas for the player who prefers not to play 34 times his minimum bet if possible, it is debatable whether any Fibonacci is any better than Martingale, or just gives you 2 extra steps and then helps you fall off the cliff more often. Therefore, no Fibonacci is recommended for the cautious player.
1-2-3-5-8-13-21-34. Go up one step with each loss, down 2 steps with each win. Note that every win pays for the two losses before it.
Example:
1. Bet $1 until you lose. Then bet $2.
2. If you win at $2, then return to step 1. If you lose, then bet $3.
3. If you win at $3, then return to step 1. If you lose, then bet $5.
4. If you win at $5, then return to step 2. If you lose, then bet $8.
5. If you win at $8, then return to step 3. If you lose, then bet $13.
6. If you win at $13, then return to step 4. If you lose, then bet $21.
7. If you win at $21, then return to step 5. If you lose, then bet $34.
8. If you win at $34, then return to step 6. If you lose, return to step 1.
Recommended variant for the high roller: instead of going down 2 steps with each win, go down 1 step with each win, until you win 2 in a row-then go down 2 steps. This variant may be as likely as any system to help you in the long run, and meanwhile you get some tremendous wins and excitement. There are numerous other variants. However the usual Fibonacci is extremely unsatisfactory for the high roller-you lose often and never win big. Whereas for the player who prefers not to play 34 times his minimum bet if possible, it is debatable whether any Fibonacci is any better than Martingale, or just gives you 2 extra steps and then helps you fall off the cliff more often. Therefore, no Fibonacci is recommended for the cautious player.
Етикети:
betting theories
COVER YOURSELF
“COVER YOURSELF” BETTING SYSTE
Bet in twos. Where possible, cover your bets so that if one loses, the other makes up for the loss. If both come in you’ve won.
Whatever your strategy, always try cover yourself. If you’re placing £1 on a bet that looks like a long shot, and although 1 in 100 times it may come in and make you a fair bit of profit, place a bet on something you’re think has a good chance of coming in that will win you that £1 you’re likely to lose on a longshot.
Let’s look at some examples :
You put £1.50 on a Lucky 15 longshot. Lucky 15s are unlikely to win every week, but when they do you could be making a large profit. So, you expect the bet to not return anything. So, place £1.50 on Liverpool to win away at Leicester at 1/1. If Liverpool win you’ve made the money you’ve written off on the speculative longshot.
You’ve got £10 to bet, and want to place it on something reasonable that has a good chance of winning. Chelsea are playing away at Middlesbrough at 6/5 and Leeds are playing away at Southampton at 11/8. Instead of putting your whole £10 on one game, don’t put all your eggs in one basket. Instead use the higher odds game as a cover, putting a smaller amount on this game. So, in this case, Leeds are your cover, and Chelsea your winner. If both win you’ve made a nice profit. Put £4.21 on Leeds at 11/8 (£10/((11/8) + 1)), which will win you exactly £10 if Leeds win, and put the rest - £5.79 on Chelsea which will win you £12.74, a small profit if only Chelsea win. If both win you’ve got £22.74 from a £10 bet. You could have placed your whole tenner on Chelsea and lost the lot on one game
Bet in twos. Where possible, cover your bets so that if one loses, the other makes up for the loss. If both come in you’ve won.
Whatever your strategy, always try cover yourself. If you’re placing £1 on a bet that looks like a long shot, and although 1 in 100 times it may come in and make you a fair bit of profit, place a bet on something you’re think has a good chance of coming in that will win you that £1 you’re likely to lose on a longshot.
Let’s look at some examples :
You put £1.50 on a Lucky 15 longshot. Lucky 15s are unlikely to win every week, but when they do you could be making a large profit. So, you expect the bet to not return anything. So, place £1.50 on Liverpool to win away at Leicester at 1/1. If Liverpool win you’ve made the money you’ve written off on the speculative longshot.
You’ve got £10 to bet, and want to place it on something reasonable that has a good chance of winning. Chelsea are playing away at Middlesbrough at 6/5 and Leeds are playing away at Southampton at 11/8. Instead of putting your whole £10 on one game, don’t put all your eggs in one basket. Instead use the higher odds game as a cover, putting a smaller amount on this game. So, in this case, Leeds are your cover, and Chelsea your winner. If both win you’ve made a nice profit. Put £4.21 on Leeds at 11/8 (£10/((11/8) + 1)), which will win you exactly £10 if Leeds win, and put the rest - £5.79 on Chelsea which will win you £12.74, a small profit if only Chelsea win. If both win you’ve got £22.74 from a £10 bet. You could have placed your whole tenner on Chelsea and lost the lot on one game
Етикети:
betting theories
TEN PERCENT
TEN PERCENT
“TEN PERCENT” BETTING SYSTEM
With each loss, go up 10% rounded to the nearest non-zero bet unit, with each win go down 10%. After you break even, then with each extra win, return halfway back to the lowest bet.
This is identical to D’Alembert except that if you hit $20, you then go up and down $2. If you hit $30, you then go up and down $3. Etc. Be sure to keep track of your total since the last time you saw $1, and reduce bets by 50% after each win that causes a total net win or that breaks even.
The Ten Percent Solution is for D’Alembert players who never want to give up, or for “cancellation” players who want to go the limit with slightly less insanity. With the Ten Percent Solution, the higher you play the more you might win-instead of being doomed to play higher and higher with less and less hope or reason.
The Ten Percent Solution can be efficient. Be warned however that this is not a low risk strategy. It is only “lower-risk play for high risk players.” The Ten Percent Solution is not for those who lack discipline.
“TEN PERCENT” BETTING SYSTEM
With each loss, go up 10% rounded to the nearest non-zero bet unit, with each win go down 10%. After you break even, then with each extra win, return halfway back to the lowest bet.
This is identical to D’Alembert except that if you hit $20, you then go up and down $2. If you hit $30, you then go up and down $3. Etc. Be sure to keep track of your total since the last time you saw $1, and reduce bets by 50% after each win that causes a total net win or that breaks even.
The Ten Percent Solution is for D’Alembert players who never want to give up, or for “cancellation” players who want to go the limit with slightly less insanity. With the Ten Percent Solution, the higher you play the more you might win-instead of being doomed to play higher and higher with less and less hope or reason.
The Ten Percent Solution can be efficient. Be warned however that this is not a low risk strategy. It is only “lower-risk play for high risk players.” The Ten Percent Solution is not for those who lack discipline.
Етикети:
betting theories
surebet
Surebet!!!!!
A SureBet is a bet, where there is such a difference in the bookmakers odds that you by betting on all outcomes in an event is sure of winning money.
How the return are calculated?
Let’s suppose a football match three outcomes, and the BEST odds:
2.80-3.50-3.10.
$1 / 2.80 = $0.36
will give you $1 in case of a home victory and get that on draw $0.29 and on away invested $0.32 to receive $1.
But
0.36 + 0.29 + 0.32 = ONLY 0.97<1 !!!!!!
For $0.97 you get back FOR SURE $1, it means $1/0.97=3.58% as return, without risk in theory.
Of course, no bookmaker offering this advantage, therefore you must be client at least three bookies. But if you can calculate this simple method, you will always have a free launch - which is a rare gift nowadays.
And finally if AH is 0:0 you get your money back when game ends up as draw.
But be aeare some bookies charge 5-10% of the stake as commission in case of refund
A SureBet is a bet, where there is such a difference in the bookmakers odds that you by betting on all outcomes in an event is sure of winning money.
How the return are calculated?
Let’s suppose a football match three outcomes, and the BEST odds:
2.80-3.50-3.10.
$1 / 2.80 = $0.36
will give you $1 in case of a home victory and get that on draw $0.29 and on away invested $0.32 to receive $1.
But
0.36 + 0.29 + 0.32 = ONLY 0.97<1 !!!!!!
For $0.97 you get back FOR SURE $1, it means $1/0.97=3.58% as return, without risk in theory.
Of course, no bookmaker offering this advantage, therefore you must be client at least three bookies. But if you can calculate this simple method, you will always have a free launch - which is a rare gift nowadays.
And finally if AH is 0:0 you get your money back when game ends up as draw.
But be aeare some bookies charge 5-10% of the stake as commission in case of refund
Етикети:
betting theories
Subscribe to:
Posts (Atom)